# KA-GNN: Kolmogorov-Arnold Graph Neural Networks for Molecular Property Prediction

### v1

@@ A.3 Sensitivity
A.4 Fourier KAN VS MLP
In contrast to the multilayer perceptron (MLP) approach, which increases the number of layers to enhance the model’s ability to fit the data, the Fourier KAN model enhances its fitting ability by adjusting the number of Fourier basis functions (controlled by the K parameter) and the option to include a bias term to decompose complex functions into simpler sine and cosine waveforms. This strategic decomposition not only improves the model’s expressiveness but also reduces the computational burden. We evaluated the fitting ability of one-layer MLP and one-layer Fourier KAN on six different function types, and the results are shown in Figure 3 and Figure 4. Our analysis shows that Fourier KAN captures the underlying properties of the function more skillfully and exhibits superior expressiveness, especially in the ability to fit periodic functions.
Figure 3: Fourier KAN fits six different functions. A. Logarithmic function $K=100$ B. Sin and Cos function $K=10$ C. Linear function $K=200$ D. Sin function $K=10$ E. Polynomial function $K=500$ F. Exponential function $K=120$
Figure 4: MLP fits six different functions. A. Logarithmic function B. Sin and Cos function C. Linear function D. Sin function E. Polynomial function F. Exponential function

### v2

@@ 3.3 Baseline Models
3.4 Hyperparameter Sensitivity Analysis
Table 6 presents an analysis of the sensitivity of various hyperparameters, including batch size, learning rate (LR), and network depth, on the performance of molecular property prediction models. The results suggest that selecting an optimal batch size and a lower learning rate tends to enhance model performance. However, increasing the number of layers does not consistently yield improvements and may even result in overfitting or diminishing returns. These findings highlight the critical role of hyperparameter tuning in achieving optimal model performance.
4 Conclusion
In this study, we propose the first non-trivial Kolmogorov-Arnold Network-based Graph Neural Networks (KA-GNNs). Our KA-GNNs utilize KAN’s unique power to optimize GNN architectures at three major levels, including node embedding, message passing, and readout. We develop Fourier series-based KAN model and provide a rigorous mathematical prove of the robust approximation capability of the Fourier KAN architecture. We compare the effects of various pre-activation function choices on the expressive power of the KAN model in terms of model performance and computational efficiency, identifying Fourier series as the optimal pre-activation function. Based on widely-used benchmark datasets for molecular property prediction, we extensively compare with existing state-of-the-art models, and find that our KA-GNNs can outperform traditional GNN models. This work not only highlights the great power of KA-GNNs in molecular property prediction but also provides a novel geometric deep learning framework for the general non-Euclidean data analysis.